how do you solve equations involing ages? for exampe:
Tom is 4 years older than Jerry.Nine years ago,Tom was 5 times as old as Jerry was then. How old is each now?
thank you

Answer by stanbon(75887)

Tom is 4 years older than Jerry.Nine years ago,Tom was 5 times as old as Jerry was then. How old is each now?
------
Equations:
T = J+4
(T-9) = 5(J-9)
----
Substitute for "T" and solve for "J":
(J+4-9) = 5J-45
----
J-5 = 5J-45
J = 5J - 40
---
4J = 40
J = 10 yrs (Jerry's age now)
----
Since T = J+4, T = 14 yrs (Tom's age now)
=============================================
Cheers,
Stan H.
==============